OpenStudy (anonymous):

A Bike rider is traveling at a speed of 15 feet per second.write n expression for the distance the rider has traveled after seconds.Make a table that records the distance for 3.0,5.8,11.1,and 14.0 seconds.

OpenStudy (anonymous):

Can any one help me??I will fan and Medall:))

OpenStudy (phi):

do you know the formula for distance ?

OpenStudy (anonymous):

Nope:))

OpenStudy (phi):

distance = rate * time or distance = speed * time

OpenStudy (anonymous):

Sorry Yeah Ik that formula..but I have to make a table and i dont have a clue what to do on that

OpenStudy (phi):

are you saying you can't make a table for time=1,2,3, and so on , and write next to it the distance ?

OpenStudy (anonymous):

I dont understand...its confusing

OpenStudy (phi):

they give time for 3.0,5.8,11.1,and 14.0 seconds.

OpenStudy (campbell_st):

well do it this way 

OpenStudy (anonymous):

so i just make a table with the 1,2,3 and so on on the left side and then the actuall distance on the right side of the table?? @phi

OpenStudy (anonymous):

Wait... @campbell_st do i now multiply them or something?

OpenStudy (phi):

I did not read the question, they want you to use time= 3.0, 5.8, 11.1, and 14.0 seconds.

OpenStudy (campbell_st):

take each time value and multiply it by 15... write the answer in the table below the time you used.

OpenStudy (phi):

but it's the same idea: time * speed (and speed is 15 ft/sec)

OpenStudy (anonymous):

@campbell_st When i find all the values..thats the distance in feet right??

OpenStudy (phi):

speed * time = distance that means if you multiply speed times time you get distance. if we look at units, speed is feet /second and time is seconds. \[ \frac{feet}{seconds} \cdot seconds \] notice that seconds/seconds "cancel" to give 1 in other words \[ \frac{feet}{\cancel{seconds}} \cdot \cancel{seconds} = feet \]